SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (June 6, 2006)

Attendees: Al Eisner, Fang Fang, Arik Kreisel, Selina Li, Rahmat Rahmat

Operations

  • Trigger recovered without much trouble at all after the power outage at IR2 on Saturday June 3rd.
  • PEP continues to deliver high luminosity (~9-10x10^33) to BaBar. We will be taking off-resonant data for about one week after this morning's MD.

DCT/DCZ

  • No TSF damage for the last 10 days.

L3

  • There was one L3 node trim during run 65428 while EMC had out-of-synch problem as well.

Data Quality and Monitoring

  • After the power outage on Saturday, the server for post processing wasn't started until Sunday morning. So, the .ps archive fastmon plots weren't made properly for runs 65384-65401 (inclusive) and they are being made by-hand afterward. Al pointed out that maybe the server should be configured in such a way to be able to automatically restart after an outage. We are not sure if Jim Hamilton can do anything about this.
  • Data quality check for Run 1-5 data for ICHEP is done.
  • Rahmat said there was no difference in TRG data with or without trickle injections. Al added that DCH found a few % efficiency lost in the region behind the inhibit window.
  • Al checked Rahmat's error calculation for the lumi ratio posted last week, and everything was fine. Rahmat will go ahead to put that in for strip charts and show us some examples next week. There is going to be an OPR release in 1-2 weeks and Rahmat will check to see if it's possible to include the lumi ratio strip charts for this release.

Trickle Injection Monitoring

  • Al showed some some trickle injection monitoring plots for run 65268 and a few runs (65289-65297) merged together. In page 2, there was a depletion band running verticl right after the inhibit window due to DAQ deadtime according to Al and Matt Weaver. Al assumes that the deatime is independent of phase and made 15 horizontal slices in the weaver planes (page 2) to produce the 1-D projection plots in the rest of the pages in these two files. By making this assumption, he can divide one histogram in one slice by the other to avoid the bias introduced by the deadtime. See HN posted after the meeting.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.